# Inverse analysis of granular flows using differentiable graph neural network simulator

### v2

@@ 4.1 Single parameter inverse
Estimating the material properties of granular media resulting in a landslide or debris flow is important for planning mitigation measures. In this problem, we aim to evaluate the friction angle $\phi$ of the granular column that produces a target runout distance $d_{\phi_{target}}$ (as outlined in fig. 1a), where $d_{\phi}$ is the distance between the flow toe and the leftmost boundary. We use AD-GNS to generate a forward simulation to estimate $d_{\phi}$ for a given initial $\phi$. We evaluate a loss as the squared error between $d_{\phi_{target}}$ and $d_{\phi}$ as $(d_{\phi}-d_{\phi_{target}})^{2}$, then use AD-GNS to compute the gradient $\nabla_{\phi}J_{\phi}$. Using GD, we update the $\phi$ to find the target $\phi$ for a desired runout $d_{\phi_{target}}$ until the squared error becomes lower than 0.0005, the optimization threshold.
@@
To test our method’s versatility, we selected four granular column collapse scenarios with varying aspect ratios and friction angles not covered in our training data. These scenarios ranged from columns with a small aspect ratio (short column, $a=0.5$) to those with a large aspect ratio (tall column, $a=2.0$).
@@
Table 2 summarizes the scenarios and results for the inversely estimated friction angles. Figure 5a shows the history of the optimization based on reverse-mode AD, and fig. 7 shows the visual progress of the optimization for all the test scenarios. We set the learning rate $\eta$ of 500 in eq. 1, and the initial guess of the friction angle starts from $30\degree$. As can be seen in fig. 5a, the optimization successfully converges to the friction angles $\phi$ that are close enough to the target values $\phi_{target}$ showing errors less than 8.94% (see table 2) in a few iterations. We also compare the predicted runout distance $d_{\phi}^{GNS}$ with the inferred $\phi$ and target runout distance $d_{\phi_{target}}$. The predictions show accurate values with a maximum error of 3.49% (see table 2). Although the inverse estimation of $\phi$ includes a small error (about $1\degree$ to $3\degree$) compared to $\phi^{target}$, the runout predictions are accurate within 3.49%. The major source of the error is the differences in runout distance computed by GNS and MPM. $d_{\phi_{target}}$ and $d_{\phi_{target}}^{GNS}$ represents the runout computed by MPM and GNS at $\phi_{target}$, respectively. Although GNS and MPM use the same friction angle, the GNS includes a small amount of error (1.12% to 5.58%) since GNS is a learned surrogate for MPM.
@@
Comparing our AD-based approach to the finite difference (FD) method reveals its stability. Figure 5b shows the optimization history using FD where the gradient is estimated through two forward evaluations at $\phi$ and $\phi+\Delta\phi$. We use $\Delta\phi=0.05$ and set the same learning rate $\eta$ as AD. Despite using the same learning rate as AD, the FD method encounters challenges. The optimization is less stable, and for the tall column with $\phi_{target}=42\degree$, it fails to identify the correct friction angle with the error of about 10%, which results in $d$ error of 16%. This error stems from FD’s reliance on the chosen $\Delta\phi$ for gradient approximation, unlike AD’s precise gradient computations. Figure 6 shows the loss history of reverse-mode AD and FD. We observe the loss trajectory of FD being trapped in the flat region between $36\degree$ to $39\degree$. This is attributed to the diminutive update term $\eta\cdot\nabla_{\phi}J_{\phi}$, a consequence of underestimating gradients due to the selected $\Delta\phi$ value (= 0.05), as opposed to the accurate gradients provided by AD. Additionally, first-order FD approximates include an error of $\mathcal{O}(\Delta\phi)$ in gradient computation.
@@
Figure 5: Optimization history of friction angle with iterations from (a) reverser-mode AD and (b) finite differentiation (FD).
@@ 4.2 Multi-parameter inverse
Figure 9 visualizes the GNS prediction of runout during optimization with different $\boldsymbol{v}$. We update the parameters by updating the ADAM with the learning rate $\eta=0.1$. The loss is defined as the mean squared error (MSE) of the coordinate values for materials points between the predicted final deposit $\boldsymbol{S}_{\boldsymbol{v}}$ and the target final deposit $\boldsymbol{S}_{\boldsymbol{v_{target}}}$ from MPM. The gray shade shows the final runout deposit produced from our scenario (fig. 8) simulated using MPM, which is our target, and the yellow dots are the GNS prediction of the final deposit at the corresponding optimization iterations. The initial velocity guesses for each layer are set to zero, as shown by the cross (×) markers in fig. 10a. As iterations progress, the velocities for each layer gradually align with $\boldsymbol{v}_{target}$, indicated by square markers in fig. 10a, and consequently, the final deposit geometry closely matches the target at iteration 29 with MSE of 3.41e-4, as shown in fig. 10b.
@@
Since GNS is a surrogate model of the forward simulator, it inherently incorporates errors; the inferred velocities might not reproduce the same final deposit geometry in the ground truth simulator, MPM. Hence, verifying if the inferred velocities are still valid in the ground truth simulator in reproducing the target final deposit is necessary.
To confirm the validity of the inferred velocities, we compare the final deposit with the inferred velocities using MPM with the target final deposit. Figure 11 shows the MPM simulation results of the final deposit from the inferred velocity at the 29th iteration and target velocity. The gray shade represents the target deposit, and the yellow dots represent the deposit from MPM using the inferred velocities. The final deposit from the inferred velocity reasonably matches our target with a minor difference at the left boundary with MSE of 8.70e-4, shown in fig. 10b. The MSE is slightly larger than the optimization error at the 29th iteration but still significantly smaller than that of the initial iteration, which is about 0.01. The comparison suggests that our method provides reasonable velocity estimations to replicate the target behavior, even though GNS had not been exposed to the discretized layered initial velocity condition during training.
Figure 11: MPM simulation results of the final deposit from the optimized velocities and target velocities.
@@ 4.3 Design of baffles to resist debris flow
We can use AD-GNS for designing engineering structures, which involves optimizing the design parameters of structural systems to achieve a specific functional outcome. We demonstrate the use of AD-GNS in the design of the debris-resisting baffles to achieve a target runout distance.
@@
Our inverse analysis aims to optimally position the baffles to halt granular flow within a predefined area (as outlined in fig. 1c) Specifically, we optimize the $z$-locations of two barriers, denoted as $\boldsymbol{z}$, to ensure that the toe of the granular jet—defined as the centroid $\boldsymbol{c}_{\boldsymbol{z}}$ of the furthest 10% of downstream particles—stops at a specific target point $\boldsymbol{c}_{target}$. In this context, $\boldsymbol{z}$ corresponds to our parameter set $\boldsymbol{\Theta}$, and $\boldsymbol{c}_{\boldsymbol{z}}$ represents our meature of runout $\boldsymbol{R}_{\boldsymbol{\Theta}}$ in fig. 2.
The initial geometry of the granular mass (see fig. 12) is cuboid-shaped with the size of 0.3 × 0.2 × 0.7 m for $x$, $y$, and $z$ coordinates, and its lower edge is at ($x$=0.2, $y$=0.125, $z$=0.25) m. The granular mass is subjected to a uniform initial $x$-velocity of 2.0 m/s. The size of the baffles is 0.15 × 0.2 × 0.15 m. The simulation domain is 2.0 × 1.0 × 1.0 m. The GNS has never seen this baffle size and the simulation domain during training, which is two times larger than the training domain (see table 1 for comparison). These differences place our inverse solver in a more challenging circumstance.
@@
Figure 13 illustrates the visual progress of the optimization. We use ADAM for the optimization with the learning rate $\eta=0.01$. The loss is defined as the MSE between $\boldsymbol{c}_{\boldsymbol{z}}$ and $\boldsymbol{c}_{\boldsymbol{z}_{target}}$. The gray broken line shows the initial geometry of the granular mass, and the yellow dots represent the final deposit after flow ceases. An orange outline highlights the flow toe, with a purple dot marking its centroid $\boldsymbol{c}_{\boldsymbol{z}}$. The black dot marks the target centroid location $\boldsymbol{c}_{\boldsymbol{z}_{target}}$. We start the initial guess of the barrier center locations at ($x$=0.675 m, $z$=0.225 m) for the lower baffle and at ($x$=0.675 m, $z$=0.925 m) for the upper baffle, as illustrated by the gray boxes in fig. 13. As iteration proceeds, the centroid of the flow toe converges to the target centroid with the baffles $z$-located at 0.341 m and 0.812 m. As depicted in fig. 14, the loss history converges to a final loss of 4.68e-6 at iteration 23.
@@
As we discussed in section 4.2, we need to validate if the inferred baffle locations are still effective in the ground truth simulator in replicating the desired outcome. Figure 15 shows the final deposit and the centroid of the flow toe at the inferred baffle locations by MPM. Although there is a noticeable difference between the MPM centroid and the target centroid, with the validation loss of 1.89e-3, this deviation is still markedly lower than the loss at the initial iteration (2.30e-2) in fig. 13. This outcome is promising considering that our scenario extrapolates beyond the training data.
Figure 15: MPM simulation results of final deposit from the inferred baffle locations and its flow toe centroid compared to the target
An essential element in our inverse analysis framework is the implementation of gradient checkpointing on GNS rollout. This technique is highly advantageous, particularly when dealing with the tracking of gradients over long timesteps in 3D scenarios, such as the one described in this section. In the absence of gradient checkpointing, we encountered significant limitations in forward steps due to GPU memory constraints, reaching the 40 GB limit within merely three steps of the rollout for the backpropagation. However, by employing gradient checkpointing and thus storing the sparse results of intermediate steps, we were able to track gradients effectively for hundreds of steps. This not only alleviated memory issues but also ensured the feasibility of our approach in larger and more complex scenarios.
@@ 4.4 Optimization efficiency
We compare the computation time per optimization iteration between our method and the baseline method (table 3). The results are reported based on the single parameter optimization scenario about tall column ($a=2.0$) with $\phi=42\degree$ shown in fig. 7d. The baseline method involves executing two forward simulations ($\boldsymbol{X}_{0},\ \boldsymbol{X}_{1},\ \ldots,\ \boldsymbol{X}_{k}$), where $k$=374 employing our ground-truth simulator, MPM, at $\phi$ and $\phi+\Delta\phi$ for estimating the gradient using FD. The computation is conducted on 56 cores of Intel Cascade Lake processors on the Texas Advanced Computing Center (TACC) Frontera. The mean computation time for the MPM forward simulation takes 8078 s with a standard deviation of 381 s, and it requires the same amount of time to conduct another forward simulation at $\phi+\Delta\phi$ to compute the gradient. Consequently, one optimization iteration requires approximately 8078$\times$2 = 16157 s, which makes the $\phi$ estimation unreasonably time-consuming. Our approach (AD-GNS) does the forward simulation only at $\phi$, and the gradient is computed based on reverse-mode AD. The computation is conducted on RTX with 16 GB memory on TACC Frontera. The forward simulation takes 33 s with the standard deviation of 2.01 s, and the gradient computation using backpropagation takes 74 s with the standard deviation of 1.36 s. Consequently, one optimization iteration requires approximately 107 s on average, outperforming our baseline case by 151 times speedup. The optimization result infers accurate $\phi$ with 6.06% of error as discussed in table 2.
We also evaluate the computation time for FD with GNS (see table 3). The gradient computation from FD with GNS requires less time than AD, since it only requires one more forward evaluation in this single parameter case. However, the optimization result from FD shows the unreasonable $\phi$ inference due to inaccurate gradient estimation, in contrast to AD which infers accurate $\phi$.
For more complex inverse problems that involve high-dimensional parameter space, such as the cases in section 4.2 and section 4.3, FD with MPM becomes almost infeasible due to the computation intensity. Although FD with GNS can return the gradient values faster than MPM, the computation time proportionally increases with the number of parameters, and the gradient values are not accurate. In contrast, our approach (AD with GNS) can still accomplish efficient and accurate gradient computation owing to the reverse-mode AD despite with increase of parameters, resulting in the successful parameter inference as we showed in section 4.2 or section 4.3.
@@
Method Forward
@@ 5 Limitations
GNS is an efficient surrogate model for forward simulation. However, generalization beyond training data may not perfectly replicate the high fidelity ground truth behaviors, resulting in some inevitable error, as shown in figs. 11 and 15. To mitigate this, expanding the training dataset to encompass a more comprehensive array of scenarios could help reduce the prediction error.

### v3

@@ 4.1 Single parameter inverse
Estimating the material properties of granular media resulting in a landslide or debris flow is important for planning mitigation measures. In this problem, we aim to evaluate the friction angle $\phi$ of the granular column that produces a target runout distance $d_{\phi_{target}}$ (as outlined in fig. 2a), where $d_{\phi}$ is the distance between the flow toe and the leftmost boundary. We use AD-GNS to generate a forward simulation to estimate $d_{\phi}$ for a given initial $\phi$. We evaluate a loss $J_{\phi}$ as the squared error between $d_{\phi_{target}}$ and $d_{\phi}$ (eq. 14), then use AD-GNS to compute the gradient $\nabla_{\phi}J_{\phi}$. Using GD, we update the $\phi$ to find the target $\phi$ for a desired runout $d_{\phi_{target}}$ until the squared error becomes lower than 0.0005, the optimization threshold.
$J_{\phi}=(d_{\phi}-d_{\phi_{target}})^{2}$ (14)
@@
To test our method’s versatility, we selected four granular column collapse scenarios with varying aspect ratios and friction angles not covered in our training data. These scenarios ranged from columns with a small aspect ratio (short column, $a=0.5$) to those with a large aspect ratio (tall column, $a=2.0$). Both cases have 3,200 particles in the simulation.
@@
Table 2 summarizes the scenarios and results for the inversely estimated friction angles. Figure 6a shows the optimization history based on reverse-mode AD, and fig. 8 shows the visual progress of the optimization for all the test scenarios. We set the learning rate $\eta$ of 500 in eq. 6, and the initial guess of the friction angle starts from $30\degree$. As can be seen in fig. 6a, the optimization successfully converges to the friction angles $\phi$ that are close enough to the target values $\phi_{target}$ showing errors less than 8.94% (see table 2) in a few iterations. We also compare the predicted runout distance $d_{\phi}^{GNS}$ with the inferred $\phi$ and target runout distance $d_{\phi_{target}}$. The predictions show accurate values with a maximum error of 3.49% (see table 2). Although the inverse estimation of $\phi$ includes a small error (about $1\degree$ to $3\degree$) compared to $\phi^{target}$, the runout predictions are accurate within 3.49%. The major source of the error is the differences in runout distance computed by GNS and MPM. $d_{\phi_{target}}$ and $d_{\phi_{target}}^{GNS}$ represents the runout computed by MPM and GNS at $\phi_{target}$, respectively. Although GNS and MPM use the same friction angle, the GNS includes a small amount of error (1.12% to 5.58%) since GNS is a learned surrogate for MPM.
@@
Comparing our AD-based approach to the finite difference (FD) method reveals its stability. Figure 6b shows the optimization history using FD where the gradient is estimated through two forward evaluations at $\phi$ and $\phi+\Delta\phi$. We use $\Delta\phi=0.05$ and set the same learning rate $\eta$ as AD. Despite using the same learning rate as AD, the FD method encounters challenges. The optimization is less stable, and for the tall column with $\phi_{target}=42\degree$, it fails to identify the correct friction angle with the error of about 10%, which results in $d$ error of 16%. This error stems from FD’s reliance on the chosen $\Delta\phi$ for gradient approximation, unlike AD’s precise gradient computations. Figure 7 shows the loss history of reverse-mode AD and FD. We observe the loss trajectory of FD being trapped in the flat region between $36\degree$ to $39\degree$. This is attributed to the diminutive update term $\eta\cdot\nabla_{\phi}J_{\phi}$, a consequence of underestimating gradients due to the selected $\Delta\phi$ value (= 0.05), as opposed to the accurate gradients provided by AD. The first-order forward FD approach has an error of $\mathcal{O}(\Delta\phi)$ in gradient computation.
@@
Figure 6: Optimization history of friction angle with iterations from (a) reverse-mode AD and (b) finite differentiation (FD).
@@ 4.2 Multi-parameter inverse
Figure 10 visualizes the GNS prediction of runout during optimization with different $\boldsymbol{v}$. We update the parameters by updating the ADAM with the learning rate $\eta=0.1$. The loss $J_{\boldsymbol{v}}$ is defined as the mean squared error (MSE) of the coordinate values for materials points between the predicted final deposit $\boldsymbol{S}_{\boldsymbol{v}}=\{\boldsymbol{p}^{i}\}_{i=1:N}$ and the target final deposit $\boldsymbol{S}_{\boldsymbol{v_{target}}}=\{\boldsymbol{p}_{target}^{i}\}_{i=1:N}$ from MPM (eq. 15). Here, $\boldsymbol{p}_{\boldsymbol{v}}^{i}$ and $\boldsymbol{p}^{i}_{\boldsymbol{v}_{target}}$ denotes the coordinates of $N$ materials points.
$J_{\boldsymbol{v}}=\frac{1}{N}\sum_{i=1}^{N}\|\boldsymbol{p}_{\boldsymbol{v}}^{i}-\boldsymbol{p}^{i}_{\boldsymbol{v}_{target}}\|^{2}$ (15)
[…]
@@
Since GNS is a surrogate model of the forward simulator, it inherently incorporates errors; the inferred velocities might not reproduce the same final deposit geometry in the ground truth simulator, MPM. Hence, verifying if the inferred velocities are still valid in the ground truth simulator when reproducing the target final deposit is necessary.
To confirm the validity of the inferred velocities, we compare the final deposit with the inferred velocities using MPM with the target final deposit. Figure 12 shows the MPM simulation results of the final deposit from the inferred velocity at the 29th iteration and target velocity. The gray shade represents the target deposit, and the red dots represent the deposit from MPM using the inferred velocities. The final deposit from the inferred velocity reasonably matches our target with a minor difference at the left boundary with an MSE of 8.70e-4, shown in fig. 11b. The MSE is slightly larger than the optimization error at the $29^{\mathrm{th}}$ iteration but still significantly smaller than that of the initial iteration, which is about 0.01. The comparison suggests that our method provides reasonable velocity estimations to replicate the target behavior, even though GNS had not been exposed to the discretized layered initial velocity condition during training.
Figure 12: MPM simulation results of the final deposit from the optimized velocities (red dots) and target velocities (gray shade), plotted together with the GNS prediction with target velocities (blue dashed line). The GNS prediction of the final deposit with optimized velocities is plotted with the yellow dashed line, which is the same as the deposit geometry from fig. 10 at the $29^{\mathrm{th}}$ iteration. The MSE at the top of the figure is for the error between the MPM with optimized velocities and the target.
@@ 4.3 Design of baffles to resist debris flow
We can use AD-GNS to design engineering structures, which involves optimizing the design parameters of structural systems to achieve a specific functional outcome. We demonstrate the use of AD-GNS in the design of the debris-resisting baffles to achieve a target runout distance.
@@
Our inverse analysis aims to optimally position the baffles to halt granular flow within a predefined area (as outlined in fig. 2c) Specifically, we optimize the $z$-locations of two barriers, denoted as $\boldsymbol{z}$, to ensure that the toe of the granular jet—defined as the centroid $\boldsymbol{c}_{\boldsymbol{z}}$ of the furthest 10% of downstream particles—stops at a specific target point $\boldsymbol{c}_{\boldsymbol{z}_{target}}$. In this context, $\boldsymbol{z}$ corresponds to our parameter set $\boldsymbol{\Theta}$, and $\boldsymbol{c}_{\boldsymbol{z}}$ represents our measure of runout $\boldsymbol{R}_{\boldsymbol{\Theta}}$ in fig. 3.
The initial geometry of the granular mass (see fig. 13) is cuboid-shaped with the size of 0.3 × 0.2 × 0.7 m for $x$, $y$, and $z$ coordinates, and its lower edge is at ($x$=0.2, $y$=0.125, $z$=0.25) m. The granular mass is subjected to a uniform initial $x$-velocity of 2.0 m/s. The size of the baffles is 0.15 × 0.2 × 0.15 m. The total number of particles is 13,371. The simulation domain is 2.0 × 1.0 × 1.0 m. The GNS has never seen this baffle size and the simulation domain during training, which is two times larger than the training domain (see table 1 for comparison). These differences place our inverse solver in a more challenging circumstance.
@@
[…]
$J_{\boldsymbol{c}_{\boldsymbol{z}}}=\|\boldsymbol{c}_{\boldsymbol{z}}-\boldsymbol{c}_{\boldsymbol{z}_{target}}\|^{2}$ (16)
The gray broken line shows the initial geometry of the granular mass, and the yellow dots represent the final deposit after flow ceases. An orange outline highlights the flow toe, with a purple dot marking its centroid $\boldsymbol{c}_{\boldsymbol{z}}$. The black dot marks the target centroid location $\boldsymbol{c}_{\boldsymbol{z}_{target}}$. We start the initial guess of the barrier center locations at ($x$=0.675 m, $z$=0.225 m) for the lower baffle and at ($x$=0.675 m, $z$=0.925 m) for the upper baffle, as illustrated by the gray boxes in fig. 14. As iteration proceeds, the centroid of the flow toe converges to the target centroid with the baffles $z$-located at 0.341 m and 0.812 m. As depicted in fig. 15, the loss history converges to a final loss of 4.68e-6 at iteration 23.
@@
As we discussed in section 4.2, we need to validate if the inferred baffle locations are still effective in the ground truth simulator in replicating the desired outcome. Figure 16 shows the final deposit and the centroid based on the optimized baffle locations using MPM. Although there is a noticeable difference between the MPM centroid and the target centroid, with the validation loss of 1.89e-3, this deviation is still markedly lower than the loss at the initial iteration (2.30e-2) in fig. 14. This outcome is promising because we extrapolate beyond the training data.
Figure 16: MPM simulation results of final deposit from the inferred baffle locations and its flow toe centroid compared to the target.
An essential element in our inverse analysis framework is the implementation of gradient checkpointing on GNS rollout. This technique is highly advantageous, particularly when tracking gradients over long timesteps in 3D scenarios, such as the one described in this section. Without gradient checkpointing, we encountered significant limitations in forward steps due to GPU memory constraints, reaching the 40 GB limit within merely three steps of the rollout for the backpropagation. However, by employing gradient checkpointing and thus storing the sparse results of intermediate steps, we could track gradients effectively for hundreds of steps. This alleviated memory issues and ensured our approach’s feasibility for larger and more complex scenarios.
@@ 4.4 Optimization efficiency
We compare the computation time between our proposed approach and the baseline Finite Difference methods (table 3). The results are reported based on the single parameter optimization scenario about the tall column ($a=2.0$) with $\phi=42\degree$ shown in fig. 8d. The baseline method involves executing two forward simulations ($\boldsymbol{X}_{0},\ \boldsymbol{X}_{1},\ \ldots,\ \boldsymbol{X}_{k}$), where $k$=374 employing our ground-truth simulator, MPM, at $\phi$ and $\phi+\Delta\phi$ for estimating the gradient using FD. The computation is conducted on 56 Intel Cascade Lake processor cores on the Texas Advanced Computing Center (TACC) Frontera. The mean computation time for the MPM forward simulation takes 8078 s with a standard deviation of 381 s, and it requires the same amount of time to conduct another forward simulation at $\phi+\Delta\phi$ to compute the gradient. Consequently, one optimization iteration requires approximately 8078$\times$2 = 16157 s, which makes the $\phi$ estimation unreasonably time-consuming. Our approach (AD-GNS) does the forward simulation only at $\phi$, and the gradient is computed based on reverse-mode AD. The computation is conducted on RTX with 16 GB memory on TACC Frontera. The forward simulation takes 33 s with the standard deviation of 2.01 s, and the gradient computation using backpropagation takes 74 s with the standard deviation of 1.36 s. Consequently, one optimization iteration requires approximately 107 s on average, outperforming our baseline case by 151 times speedup. The optimization result infers accurate $\phi$ with 6.06% of error as discussed in table 2.
We also evaluate the computation time for FD with GNS (see table 3). The gradient computation from FD with GNS requires less time than AD since it only requires one more forward evaluation in this single parameter case. However, the optimization result from FD shows the unreasonable $\phi$ inference due to inaccurate gradient estimation, in contrast to AD, which infers accurate $\phi$.
For more complex inverse problems that involve high-dimensional parameter space, such as the cases in section 4.2 and section 4.3, FD with MPM becomes almost infeasible due to the computation intensity. Although FD with GNS can return the gradient values faster than MPM, the computation time proportionally increases with the number of parameters, and the gradient values are inaccurate. In contrast, our approach (AD with GNS) can still accomplish efficient and accurate gradient computation owing to the reverse-mode AD despite the increase in the number of parameters, resulting in the successful parameter inference as we showed in section 4.2 or section 4.3.
@@
{tblr}
cells = c, cell11 = r=2, cell12 = c=2, cell14 = c=2, cell16 = r=2, cell17 = r=2, cell18 = r=2, cell34 = c=2, cell54 = c=2, hline1,6 = -0.08em, hline2 = 2-5, hline3 = -, Method & Forward
@@ 5 Limitations
GNS is an efficient surrogate model for forward simulation. However, generalization beyond training data may not perfectly replicate the high fidelity ground truth behaviors, resulting in some inevitable error, as shown in figs. 12 and 16. To mitigate this, expanding the training dataset to encompass a more comprehensive array of scenarios could help reduce the prediction error.
While gradient checkpointing offers a workaround for the high memory demands associated with reverse-mode AD in GNN, tracking the gradient over the GNN with large graphs is still expensive. We evaluated peak memory usage during backpropagation across up to 400 forward computation steps with increasing particles, as depicted in fig. 17. At around 20K particles, the backpropagation process exceeds the memory capacity of our device (16 GB, RTX node at Frontera, TACC), failing. Introducing checkpoints more frequently can reduce memory consumption; however, this strategy requires additional computations for the forward pass, leading to a trade-off between memory efficiency and computational overhead. Another strategy to alleviate memory constraints involves partitioning the graph (Çatalyürek et al., 2023) and employing multi-GPU processing. Our GNS implementation (Kumar and Vantassel, 2022) supports efficient multi-GPU processing. However, effectively partitioning large graphs and facilitating communication across these partitions in a multi-GPU environment presents challenges and is the subject of active research.
Figure 17: Peak memory usage during backpropagation across up to 400 forward computation steps in relation to an increasing number of particles.